# MATHWELL: Generating Educational Math Word Problems at Scale

### v3

@@ Abstract
Math word problems are critical K-8 educational tools, but writing them is time-consuming and requires domain expertise. We suggest that language models can support K-8 math education by automatically generating problems at scale. To be educational, generated problems must be 1) solvable, 2) accurate, and 3) appropriate. Existing datasets are unlabeled for these criteria, making them ill-suited for training problem generators. We introduce MATHWELL, a Llama-2 (70B) model iteratively finetuned to generate K-8 math word problems using data from expert annotation. Using MATHWELL, we generate the largest English word problem dataset with Program of Thought (PoT) rationales to date, containing 20,490 problems. 3,484 are scored by domain experts who find MATHWELL has a 40% higher share of problems that have executable solutions and meet all criteria than alternatives, with 74% of its problems with executable solutions being solvable, accurate, and appropriate. We release our model, data, and annotations.111https://github.com/bryanchrist/MATHWELL
@@ 1 Introduction
Math word problems are a critical part of assessing student learning in K-12 education Daroczy et al. (2015); Pearce et al. (2013); Schwartz (2023); Verschaffel et al. (2020). Customizing word problems to student interests can further increase learning, interest in math, and test performance Bernacki and Walkington (2018); Walkington (2013). However, with many students and responsibilities, teachers rarely have time to customize questions for their students. We aim to automatically generate customized math word problems and answers at scale to facilitate fast, interest-guided, math education.
@@
Math word problems include natural language questions paired with numerical solutions, as illustrated in Figure 1. To be effective educational tools for grade-school students, the problems must be solvable, appropriate, and have accurate solutions. While most public word problem resources are manually curated and, therefore, quite limited (e.g., K5Learning ; TPT ; VDOE ), there have been recent efforts to generate math word problems using natural language processing Jiao et al. (2023); Koncel-Kedziorski et al. (2016); Niyarepola et al. (2022); Norberg et al. (2023); Qin et al. (2023, 2024); Wang et al. (2021); Zhou and Huang (2019); Zhou et al. (2023); Zong and Krishnamachari (2023). Closest to our work, some recent efforts explore math word problem generation using Large Language Models (LLMs) Niyarepola et al. (2022); Zong and Krishnamachari (2023). However, to the best of our knowledge, existing methods have one or both of two important limitations. First, many only generate questions (e.g., Niyarepola et al. 2022; Zong and Krishnamachari 2023). To be educationally useful, students and teachers must also have access to answers. Further, it is hard to verify if questions are solvable and appropriate without answers. Second, all existing methods are context-dependent, simply rephrasing or creating new output based on input problems or equations. In practice, curating and choosing specific word problems or equations beforehand is laborious and difficult Roche (2013).
To address the aforementioned limitations, we propose context-free educational math word problem generation. Successful solutions must return questions and answers together and require no hand-picked equations or exemplar word problems. For context-free word problem generators to be effective educational tools, we propose three criteria to evaluate their outputs. 1) Each problem must be solvable. 2) Each question’s corresponding answer must be accurate. 3) Each problem must be appropriate, where the question/answer should make logical sense, the math topic should be familiar to a student, and the question’s context should be appropriate for a young learner.
@@
We propose Math Assessment Through High-interest Word Problems Extracted from LLMs (MATHWELL), the first context-free educational math word problem generator. MATHWELL is a Llama-2 (70B) model finetuned on popular math question-answering (QA) datasets used by recent math LLMs Azerbayev et al. (2023); Luo et al. (2023); Yue et al. (2023). However, these datasets are unannotated for solvability, accuracy, and appropriateness and are not designed to train word problem generators. We therefore conduct multiple rounds of annotation by real expert educators, which we use to progressively finetune MATHWELL. With these annotations, our resultant model’s outputs are better targeted for grade-school math education than outputs from alternative general-purpose math LLMs. Further, we train MATHWELL to generate Program of Thought (PoT) solutions to word problems, which recent methods show improves accuracy on open-ended questions Azerbayev et al. (2023); Gao et al. (2023); Yue et al. (2023).
@@
We use MATHWELL to generate Synthetic Grade School Math (SGSM), a dataset of 20,490 word problems. This is the largest public set of English math word problems with PoT rationales. We acquire annotations from domain experts for 3,484 model samples during and after training to score the quality of our data, resulting in a SGSM subset with 2,093 samples with gold labels for solvability, accuracy, and appropriateness. We find that MATHWELL is effective at generating custom math word problems without context: 74% of its output problems with executable code are simultaneously solvable, accurate, and appropriate and nearly all successfully include a specified topic. These outputs are similar to popular human-written math QA datasets. MATHWELL has a nearly 20% higher share of outputs with executable code that meet our three evaluation criteria than alternative LLMs. Further, we find that SGSM has a more age appropriate reading level than other datasets. We also show that our human annotations can be used to train text classifiers to automatically score MATHWELL outputs.
@@
We release MATHWELL, a context-free word problem generator, and show it has a 43.4% higher share of outputs that meet all evaluation criteria and have executable code than the next best open-source model and is the only model that generates questions that are all at an appropriate reading level for K-8 students.
We release SGSM and show it is high quality through automatic evaluation. SGSM is the only existing dataset designed to train context-free word problem generators, is the largest English math grade school dataset with PoT solutions, and contains more readable questions than existing grade school math datasets.
@@ 2 Related Work
There are several popular datasets for training LLMs to answer grade school math questions in English, including GSM8K Cobbe et al. (2021), NumGLUE Mishra et al. (2022b), GSM-Hard Gao et al. (2023), ASDIV Miao et al. (2020) and SVAMP Patel et al. (2021). Recent work has also developed new grade school math datasets or different rationales for existing datasets, including those that are human written Kim et al. (2023) and synthetic Chern et al. (2023); Mitra et al. (2024); Shi et al. (2023); Yu et al. (2023); Yue et al. (2023). However, because existing datasets were created to train LLMs to answer, rather than generate, grade school math problems, they do not possess the characteristics we consider critical to training LLMs to generate solvable, accurate, and appropriate question/answer pairs. Specifically, while some datasets include PoT solutions (e.g., GSM-Hard, MathInstruct GSM8K, and NumGLUE), only GSM-Hard contains Python function solutions, which we find increases the likelihood that a question/answer pair contains executable code. Because GSM-Hard contains questions with inflated number values, though, it is not appropriate for training a question generator for K-8 students. Due to the limitations of existing datasets, we create and release SGSM, a SGSM subset verified by human annotation that contains Python function solutions and appropriate number ranges for K-8 students.
@@
Other work explores automatically generating math word problems but takes a more limited approach than the present study. The closest related work is Zong and Krishnamachari (2023), who assess GPT-3’s ability to generate math word problems. Their method is not context-free, however, as they use a reference problem to guide generation. In addition, they only generate word problems, not solutions. Niyarepola et al. (2022) train LLM word problem generators, but they guide generation by providing the beginning of a word problem and do not generate solutions. Most other existing works use LLMs or deep neural networks to generate word problems based on pre-specified equations, provide additional questions based on reference word problems, or re-write existing word problems Jiao et al. (2023); Koncel-Kedziorski et al. (2016); Norberg et al. (2023); Qin et al. (2023, 2024); Wang et al. (2021); Zhou and Huang (2019); Zhou et al. (2023), each of which are more limited than context-free generation by restricting the range of possible outputs. These approaches also require additional input from users, which may not be feasible for teachers or students who wish to create customized word problems. As a result, we do not compare MATHWELL against these models/methods. Our study is the first to evaluate the capability of leading open-source LLMs for the task of math word problem generation.
@@ 3 Methods
We propose MATHWELL, a finetuned language model that generates word problems that align with topics K-8 students are interested in. MATHWELL can be prompted to generate word problems, where each problem is a question/answer pair. Then, we use this language model to generate a large synthetic dataset, which we evaluate with human domain experts. We illustrate our model finetuning and dataset generation process in Figure 2.
There are three critical criteria for model-generated question/answer pairs: 1) Solvability indicates questions are possible to solve and have one correct answer Jiao et al. (2023); Koncel-Kedziorski et al. (2016); Qin et al. (2023). 2) Accuracy measures if the generated answer is correct. 3) Appropriateness indicates that questions/answers make logical sense, do not have grammatical errors or conflicting information, and are appropriate for K-8 students mathematically and contextually. A secondary characteristic we consider for model comparisons is topic specificity, or if the question incorporates the topic specified in the prompt.
@@ 3.1 Training MATHWELL
MATHWELL is a finetuned Llama-2 (70B) Touvron et al. (2023) model that generates question/answer pairs. After initial finetuning using public data, we generate synthetic data and acquire expert annotations for solvability, accuracy, and appropriateness. We then further finetune on the best samples to ensure high-quality outputs. We generate Program of Thought (PoT) solutions, which recent works have shown improves accuracy Azerbayev et al. (2023); Gao et al. (2023); Yue et al. (2023).
Initial Finetuning
We re-format MathInstruct GSM8K Yue et al. (2023) into an Alpaca-style Taori et al. (2023) instruction dataset of question/answer pairs with a standard prompt asking the model to generate a grade school math word problem. To the best of our knowledge, MathInstruct GSM8K is the only dataset with Python code solutions and questions that are all appropriate for K-8 students Yue et al. (2023). We also include MathInstruct MATH and TheoremQA PoT Yue et al. (2023) question/answer pairs with a standard prompt asking the model to generate a challenge math problem to expose the model to more complex code and further promote mathematical reasoning. We then finetune Llama-2 (70B) on this dataset of 25,926 rows using QLoRA Dettmers et al. (2023) for 4,250 steps. We use two A100 GPUs, a learning rate of 1e-4, LoRA modules at every model layer, per-device batch size of 1, and a 3% warm-up ratio.
Secondary Finetuning on Expert-annotated Data
Inspired by recent works that iteratively finetune LLMs Guo et al. (2024); Wang et al. (2024), including OpenAI’s FeedME approach Zheng et al. (2023), we next aim to evaluate and improve how well MATHWELL meets our three main criteria: solvability, accuracy, and appropriateness. However, these properties are unannotated in prior works. Therefore, we few-shot prompt the finetuned model and generate 3,234 problems. We then acquire annotations for solvability, accuracy, and appropriateness from volunteers with previous K-12 teaching experience/training. Further details about the annotation process and annotator agreement are in Appendix G. These annotations identified 1,906 problems that are simultaneously solvable, accurate, and appropriate, or meet all criteria (MaC). We then finetune MATHWELL on these MaC outputs for 1,250 more steps.
3.2 Generating a Dataset with MATHWELL
@@
We next prompt MATHWELL to generate a large dataset, which we call Synthetic Grade School Math (SGSM). SGSM contains 20,490 question/answer pairs generated by MATHWELL. SGSM has two subsets: SGSM, containing 2,093 MaC question/answer pairs, and SGSM, containing 18,397 question/answer pairs that have executable code but are not labeled for solvability, accuracy, and appropriateness. SGSM is the largest English grade school PoT math QA dataset. To support research in automatically scoring model-generated math word problems, we release our annotated data. To the best of our knowledge, this is the only human annotated model-generated math word problem dataset (see Appendix B for details).
Model | Solv. | Acc. | App. | MaC | Top. Spec. | EC | EC/MaC |
LLEMMA | 48.8 (3.17) | 63.9 (4.37) | 41.8 (4.48) | 15.2 (2.28) | 94.8 (1.41) | 24.3 (0.70) | 3.70 (0.55) |
MAmmoTH | 86.8 (2.15) | 94.9 (1.49) | 67.7 (3.18) | 56.8 (3.14) | 97.6 (0.97) | 6.90 (0.36) | 3.91 (0.22) |
Llama-2 | 84.0 (2.32) | 89.5 (2.12) | 81.0 (2.72) | 62.4 (3.07) | 99.2 (0.56) | 55.4 (0.98) | 34.6 (1.70) |
MATHWELL | 89.2 (1.97) | 96.9 (1.17) | 86.5 (2.29) | 74.8* (2.75) | 99.6 (0.40) | 66.4* (1.00) | 49.6* (1.83) |
@@
Table 1 shows the advantages of SGSM over existing datasets. First, it is the largest available PoT dataset, which may enable further training of task-specific models. Second, SGSM is the only dataset with both Program of Thought (PoT) solutions written as Python functions and questions that assess mathematical concepts appropriate for K-8 students, which we find critical for training context-free word problem generators (see Appendix A). Based on the experiments reported in Section 4, we believe SGSM is also predominately comprised of questions that are mathematically appropriate for K-8 students. Third, SGSM and its subsets have the lowest average Flesch-Kincaid grade level (FKGL) Flesch (1948); Kincaid et al. (1975) and comparable New Dale-Chall (NDC) readability score Chall and Dale (1995), so the questions may be more appropriate for those who struggle to read.
SGSM is similar to existing datasets on other metrics commonly used to evaluate math datasets and/or model-generated word problems (e.g., Jiao et al. 2023; Niyarepola et al. 2022; Norberg et al. 2023; Zhou and Huang 2019; Zhou et al. 2023). Specifically, as shown in Table 1, while SGSM has a shorter average token length than MathInstruct GSM8K Yue et al. (2023), its token length is longer than ASDIV Miao et al. (2020) and SVAMP Patel et al. (2021), two other grade school math datasets. This suggests that SGSM’s question complexity, as reflected in average question length, should be similar to existing datasets. SGSM’s BERTScore is close to those of existing datasets, suggesting the questions are similar. Taken together, these findings suggest that SGSM has important advantages over existing datasets and is comparable to human-written data.

### v4

# MATHWELL: Generating Age-Appropriate Educational Math Word Problems

@@ Abstract
Math word problems are critical K-8 educational tools, but writing them is time-consuming and requires domain expertise. We suggest that language models can support K-8 math education by automatically generating problems. To be educational, generated problems must be 1) solvable, 2) accurate, and 3) appropriate. Existing datasets are unlabeled for these criteria, making them ill-suited for training problem generators. To address this gap, we use domain expert annotation to curate a high-quality synthetic training dataset for this task. We show the value of this data by using it to iteratively finetune Llama-2 (70B) to create MATHWELL, a K-8 word problem generator. Domain experts find MATHWELL has a 40% higher share of problems that have executable solutions and meet all criteria than existing open-source models, with 74% of its problems with executable solutions being solvable, accurate, and appropriate. MATHWELL achieves 94.9% of GPT-4 Turbo’s performance on this task while outputting problems written at a more appropriate reading level for K-8 students. MATHWELL’s performance despite being trained by finetuning only highlights the quality of our synthetic data for training age-appropriate word problem generators. We release our model, data, and annotations.111https://github.com/bryanchrist/MATHWELL
@@ 1 Introduction
Math word problems are critical for assessing learning in K-12 education Daroczy et al. (2015); Pearce et al. (2013); Schwartz (2023); Verschaffel et al. (2020). Customizing word problems to student interests can further increase learning, interest in math, and test performance Bernacki and Walkington (2018); Walkington (2013); Walkington and Bernacki (2019). However, with many students and responsibilities, teachers rarely have time to customize questions for each student. We aim to automatically generate customized math word problems and answers to promote interest-guided math education.
@@
Math word problems include natural language questions paired with numerical solutions, as illustrated in Figure 1. To be effective educational tools for grade-school students, the problems must be solvable, appropriate, and have accurate solutions. While most public word problem resources are manually curated and limited (e.g., K5Learning ; TPT ; VDOE ), there have been recent efforts to generate these problems using natural language processing Jiao et al. (2023); Koncel-Kedziorski et al. (2016); Niyarepola et al. (2022); Norberg et al. (2023); Qin et al. (2023, 2024); Wang et al. (2021); Zhou and Huang (2019); Zhou et al. (2023); Zong and Krishnamachari (2023). Closest to our work, some recent efforts explore math word problem generation using Large Language Models (LLMs) Niyarepola et al. (2022); Zong and Krishnamachari (2023).
However, to the best of our knowledge, existing methods have one or both of two important limitations.
First, many only generate questions (e.g., Niyarepola et al. 2022; Zong and Krishnamachari 2023).
To be educationally useful, students and teachers must also have access to answers.
Further, it is hard to verify if questions are solvable and appropriate without answers.
Second, all existing methods are context-dependent, simply rephrasing or creating new output based on input problems or equations.
In practice, curating and choosing specific word problems or equations beforehand is laborious and difficult Roche (2013), and there are more equations and word problem templates than can reasonably be pre-specified. There are also contexts for which having increased flexibility over question structure and underlying equations may be preferable, such as preparing students for problems that mix/combine operations or allowing students to practice solving problems on their own without knowing the equation the problems will follow before generation. This flexibility can ensure students master different types of word problems rather than memorize operations associated with specific phrasing/template equations, which harms generalizability and limits overall learning Ben-Zeev and Star (2001); Chang et al. (2003).
To address the aforementioned limitations, we propose context-free age-appropriate, educational math word problem generation.
Successful solutions must return questions and answers together and require no hand-picked equations or exemplar word problems.
For context-free word problem generators to be useful educational tools, we propose three criteria to evaluate their outputs.
1) Each problem must be solvable.
2) Each question’s corresponding answer must be accurate.
3) Each problem must be appropriate, where the question/answer must make logical sense, the math topic must be familiar to a student, and the question’s content must be appropriate for a young learner.
@@
To explore context-free age-appropriate, educational math word problem generation, we introduce MATHWELL, the first such generator trained through finetuning Llama-2 (70B) in two stages. The first stage uses popular math question-answering (QA) datasets used in recent work Azerbayev et al. (2023); Luo et al. (2023); Yue et al. (2023). However, these datasets are unannotated for solvability, accuracy, and appropriateness and are not designed to train word problem generators, making them ill-suited for this task. In turn, we use our first-stage model to generate high-quality synthetic training data for age-appropriate, educational word problem generators verified by human experts. To demonstrate the data’s quality and improve our model, we use it to conduct a second stage of finetuning. We show MATHWELL’s outputs are better targeted for grade school math education than those from existing open-source LLMs.
@@
We use MATHWELL to generate Synthetic Grade School Math (SGSM), containing 20,490 word problems. We acquire annotations from domain experts for 3,484 MATHWELL outputs during and after training to score the quality of our data, resulting in a SGSM subset with 2,093 samples with gold labels for solvability, accuracy, and appropriateness. We find that MATHWELL is effective at generating custom math word problems without context: 74% of its output problems with executable code are simultaneously solvable, accurate, and appropriate and nearly all successfully include a specified topic. These outputs are similar to popular human-written math QA datasets. MATHWELL has a nearly 20% higher share of outputs with executable code that meet our three evaluation criteria than existing open-source LLMs, outperforming GPT-3.5 Turbo and matching 94.9% of GPT-4 Turbo’s performance while also generating problems with a more appropriate reading level. Further, we find that SGSM has a more age-appropriate reading level than other datasets.
@@
We release SGSM, the first dataset designed to train age-appropriate, educational context-free grade school math word problem generators, and use automatic evaluation to show it is both high quality and contains more readable questions than existing datasets.
We further evaluate SGSM by using it to finetune MATHWELL, a context-free word problem generator that has a 43.4% higher share of outputs that meet all three criteria and have executable code than the next best open-source model and nearly eliminates the gap between open and closed source models on our three criteria. MATHWELL is the only model that consistently outputs questions at an appropriate reading level for K-8 students.
@@ 2 Related Work
There are many datasets for training and evaluating LLMs on grade school math reasoning. Several popular datasets include GSM8K Cobbe et al. (2021), NumGLUE Mishra et al. (2022), GSM-Hard Gao et al. (2023), ASDIV Miao et al. (2020) and SVAMP Patel et al. (2021). Recent work has also developed new grade school math evaluation datasets or large training datasets with different solution rationales for existing datasets, particularly GSM8K, including both human written Kim et al. (2023); Mishra et al. (2023) and synthetic Mitra et al. (2024); Shi et al. (2023); Toshniwal et al. (2024); Yu et al. (2023); Yuan et al. (2023); Yue et al. (2023) data.
However, because existing datasets are designed primarily for training and evaluating grade school math reasoning, they are not aligned with training age-appropriate, educational grade school math word problem generators. Specifically, grade school math word problem training data for age-appropriate, educational generators must contain high-quality grammar, be written at an appropriate mathematical difficulty and reading level for K-8 students, and include questions similar to those students would encounter in the classroom. Further, to encourage effective solution generation, training data must contain PoT solutions written as Python functions (see Appendix B for details).
We consider high-quality grammar and similarity to word problems students encounter in the classroom baseline criteria any potentially relevant training data must possess. Existing datasets that contain these baseline characteristics and are aligned with one or more of the other three criteria are GSM-Hard Gao et al. (2023), MathInstruct GSM8K Yue et al. (2023), ASDIV Miao et al. (2020), and SVAMP Patel et al. (2021), although none of these datasets contain all three criteria. For example, while some datasets include PoT solutions (e.g., GSM-Hard and MathInstruct GSM8K), only GSM-Hard contains Python function solutions. Because GSM-Hard contains questions with inflated number values, though, it is not appropriate for training a question generator for K-8 students.
@@
Other work explores automatically generating math word problems but is not context-free and, therefore, constraints the range of possible outputs more than the present study. Closest to our work is Zong and Krishnamachari (2023), who assess GPT-3’s ability to generate math word problems. Their method is not context-free, however, as they use a reference problem to guide generation. Further, they only generate word problems, not solutions. Niyarepola et al. (2022) train LLM word problem generators, but they guide generation with the beginning of a word problem and do not generate solutions.
Most other works use LLMs or deep neural networks to generate word problems based on pre-specified equations, provide additional questions based on reference problems, or re-write existing problems Jiao et al. (2023); Koncel-Kedziorski et al. (2016); Norberg et al. (2023); Qin et al. (2023, 2024); Wang et al. (2021); Zhou and Huang (2019); Zhou et al. (2023), each of which restricts the range of possible outputs. These approaches are not context-free and require additional input from users, which may be infeasible for teachers or students who wish to create customized word problems, making them incomparable to our work.
@@ 3 Methods
We propose MATHWELL, a finetuned LLM that generates age-appropriate word problems aligned with topics K-8 students are interested in and comprised of question/answer pairs.
We use MATHWELL to generate a large synthetic dataset designed to train age-appropriate, educational grade school word problem generators, which we evaluate with human domain experts.
Figure 2 illustrates our model finetuning and data generation process.
There are three critical criteria for model-generated question/answer pairs:
1) Solvability indicates questions are possible to solve and have one correct answer Jiao et al. (2023); Koncel-Kedziorski et al. (2016); Qin et al. (2023).
2) Accuracy measures if the generated answer is correct.
3) Appropriateness indicates that questions/answers make logical sense, do not have grammatical errors or conflicting information, and are appropriate for K-8 students mathematically and contextually.
A secondary characteristic we consider for model comparisons is topic specificity, or if the question incorporates the topic specified in the prompt.
@@ 3.1 Training MATHWELL
MATHWELL is a finetuned Llama-2 (70B) Touvron et al. (2023) model that generates question/answer pairs with Program of Thought (PoT) solutions, which recent works have shown improves accuracy Azerbayev et al. (2023); Gao et al. (2023); Yue et al. (2023). After initial finetuning using public data, namely MathInstruct GSM8K Yue et al. (2023), we prompt our finetuned model to generate synthetic grade school math word problems and acquire expert annotations for solvability, accuracy, and appropriateness. All annotators had K-12 teaching experience. Further details about the annotation process and annotator agreement are in Appendix L. Our annotators identified 1,906 problems that are simultaneously solvable, accurate, and appropriate, or meet all criteria (MaC). We then finetune MATHWELL further on these MaC outputs to improve performance and demonstrate the value of our synthetic data (see Appendix E for results from each finetuning stage). We conduct all finetuning using QLoRA Dettmers et al. (2023). See Appendix A for more finetuning details.
3.2 Generating Synthetic Grade School Math
@@
After the second stage of finetuning, we use MATHWELL to continue generating synthetic word problems, resulting in a large dataset we call Synthetic Grade School Math (SGSM) that combines the samples generated after each finetuning stage. In total, SGSM contains 20,490 question/answer pairs generated by MATHWELL. SGSM has two subsets: SGSM, containing 2,093 MaC question/answer pairs verified by human experts, and SGSM, containing 18,397 question/answer pairs that have executable code but are not labeled for solvability, accuracy, and appropriateness. To support research in automatically scoring model-generated math word problems, we release our annotated data. To the best of our knowledge, this is the only human annotated model-generated math word problem dataset (see Appendix C for details on our annotated data).
| Model | Solv. | Acc. | App. | MaC | Top. Spec. | EC | EC/MaC | |
| API | GPT-4 Turbo | 94.8 (1.41) | 95.8 (1.31) | 84.4 (2.36) | 78.8 (2.59) | 99.2 (0.56) | 66.8 (2.41) | 52.6 (1.73) |
| GPT-3.5 Turbo | 88.0 (2.06) | 89.5 (2.07) | 75.5 (2.91) | 62.8 (3.06) | 98.8 (0.69) | 97.5 (0.32) | 61.3 (2.99) | |
| Public | LLEMMA | 48.8 (3.17) | 63.9 (4.37) | 41.8 (4.48) | 15.2 (2.28) | 94.8 (1.41) | 24.3 (0.70) | 3.70 (0.55) |
| MAmmoTH | 86.8 (2.15) | 94.9 (1.49) | 67.7 (3.18) | 56.8 (3.14) | 97.6 (0.97) | 6.90 (0.36) | 3.91 (0.22) | |
| Llama-2 | 84.0 (2.32) | 89.5 (2.12) | 81.0 (2.72) | 62.4 (3.07) | 99.2 (0.56) | 55.4 (0.98) | 34.6 (1.70) | |
| MATHWELL (Ours) | 89.2 (1.97) | 96.9 (1.17) | 86.5 (2.29) | 74.8* (2.75) | 99.6 (0.40) | 66.4* (1.00) | 49.6* (1.83) |
@@
Table 1 shows the advantages of SGSM compared to other datasets that have grammatically-correct questions similar to those students encounter in the classroom and are PoT, have Python function solutions, and/or are written at an appropriate difficulty for K-8 students. First, SGSM is larger than other similar PoT datasets. Second, SGSM is the only dataset with both Program of Thought (PoT) solutions written as Python functions and questions that are mathematically appropriate for K-8 students, which we find critical for training context-free word problem generators (see Appendix B). Based on the experiments reported in Section 4, we believe SGSM is also predominately comprised of questions that are mathematically appropriate for K-8 students and will enable future research in evaluating synthetic educational math word problems. Third, SGSM has the lowest average Flesch-Kincaid grade level (FKGL) Flesch (1948); Kincaid et al. (1975) and comparable New Dale-Chall (NDC) readability score Chall and Dale (1995), so the questions may be more appropriate for those who struggle to read.
SGSM is similar to existing datasets on other metrics commonly used to evaluate math datasets and/or model-generated word problems (e.g., Jiao et al. 2023; Niyarepola et al. 2022; Norberg et al. 2023; Zhou and Huang 2019; Zhou et al. 2023). Specifically, while SGSM has a shorter average token length than MathInstruct GSM8K Yue et al. (2023), its token length is longer than ASDIV Miao et al. (2020) and SVAMP Patel et al. (2021), two other grade school math datasets. This suggests that SGSM’s question complexity, as reflected in average question length, should be similar to existing datasets. SGSM’s BERTScore is close to those of existing datasets, suggesting the questions are similar. Taken together, these findings suggest that SGSM has important advantages over existing datasets and is comparable to human-written data.